A horizontal rope is tied to a 56.0kg box on frictionless ice. What is the tension in the rope if:

a) The box is at rest ?

b) The box moves at a steady

vx = 4.80m/s ?

c)
The box

vx = 4.80m/s and ax = 4.60m/s2 ?
Answer #1

a) zero


b) zero

C) T = m*ax = 257.6 N
Answer #2

(a) When the box is in rest, net horizontal force = 0,

so, T = 0 (since no other force is there)


(b) Also, when the box moves with constant velocity, net horizontal force = 0,

so, T = 0


(c) Net force in x direction = m * a

so, T = 56 * 4.60 Kg m / s^2 = 257.6 N
Answer #3

a)when the box is at rest the tension is

T = m x g

where m = 56.0 kg and g = 9.8 m/s^2

b)when the box moves at a steady velocity vx = 4.80 m/s the tension is

T1 = m x g - m x a

where a = (vx/t),t is time

c)when the acceleration of box is ax = 4.60 m/^2 then the tension is

T2 = m x g - m x a1

where a1 = 4.60 m/s^2